OpenStudy (anonymous):

the integral of e^6x/e^6x+9

OpenStudy (anonymous):

put z=denominator of d integrand

OpenStudy (anonymous):

then dz=6 e^(6x) dx

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

yeah but then what?

OpenStudy (anonymous):

I think I read it wrong not sure you can do long division.. but you can use u substitution twice

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ e^{6x}}{ e^{6x} +9}\]

OpenStudy (anonymous):

first let u = 6x... then do substitution again and use v=e^u+9

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ e ^{6x} }{ e ^{6x}+9 }\] Let: \[u=e ^{6x}+9\] Therefore: \[du = 6e ^{6x} dx\] Just divide the 6 out and put it outside the integral. The integral is now: \[\frac{ 1 }{ 6 } \int\limits_{}^{} \frac{ 1 }{ u } du = \frac{ 1 }{ 6 }\ln(u) + c\] Simply replace u with the function we originally substituted out: \[=\frac{ 1 }{ 6 }\ln(e ^{6x}+9)+c\]

OpenStudy (anonymous):

nj that was perfect. thanks for the help!

OpenStudy (anonymous):

Not a problem :D

OpenStudy (anonymous):

NJ could you help me with another please?

OpenStudy (anonymous):

Certainly :) Just post it for me to see!